# QMetro++ - Python optimization package for large scale quantum metrology with customized strategy structures

### v1

@@ Abstract
QMetro++ is a Python package containing a set of tools dedicated to identifying optimal estimation protocols that maximize quantum Fisher information (QFI). Optimization can be performed for an arbitrary arrangement of input states, parameter encoding channels, noise correlations, control operations and measurements. The use of tensor networks and iterative see-saw algorithm allows for an efficient optimization even in the regime of large number of channel uses ($N\approx 100$). Additionally, the package comes with an implementation of the recently developed methods for computing fundamental upper bounds on QFI, which serve as benchmarks of optimality of the outcomes of numerical optimization. All functionalities are wrapped up in a user-friendly interface which enables defining strategies at various levels of detail.
@@ 1 Introduction
(B) parallel small $N$ mop_parallel_qfi, iss_parallel_qfi 3.3
large $N$ iss_tnet_parallel_qfi
@@
small $N$ mop_parallel_qfi, iss_parallel_qfi 3.4 3.5
(C) large $N$ iss_tnet_adaptive_qfi
@@
Figure 2: File structure of the package. Classes from tensors.py file are displayed in more detail in Fig. 10.
@@
Quantum metrology [1, 2, 3, 4] is now rightfully regarded as one of the main pillars of rapidly developing field of quantum technologies [5]. It focuses on the ways to exploit delicate quantum features of light and matter to enhance the sensitivity of practical measuring devices. Apart from numerous experimental breakthroughs in the field, including the use of squeezed states of light in gravitational wave detectors [6, 7], quantum enhanced magnetometry [8], gravitometry [9], atomic clocks [10] and many others, the field has also enjoyed significant theoretical progress with numerous analytical and numerical methods being developed, that help to design better metrological protocols as well as understand the fundamental limitations of quantum-enhanced protocols in the presence of decoherence [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
The purpose of this package is to make these advanced and powerful numerical methods accessible to the community. We realize that despite the fact that the core elements of the methods have been discussed in the literature for a couple of years already, they are still not widely used in the field due to their somewhat steep ‘initial learning curve’ and the fact that there is no single publication that provides a unified presentation of all of them, not to mention a numerical package. Moreover, we significantly expand the applicability of the latest tensor network-based methods to include general protocol structures, which allows to deal with e.g. collisional metrological schemes.
@@
Not so rarely, we have come across research conducted in the field, where the problem of optimization of quantum metrological protocols has been approached using general-purpose optimization tools (including even the latest AI methods), which in the end did not result in a particular efficient numerical algorithm, the results of numerical optimization had not optimality guarantee and were not benchmarked with known fundamental precision bounds [citations intentionally hidden here!].
@@
This package adopts the frequentist approach to quantum metrology and focuses on quantum Fisher information (QFI) as the figure-of-merit to be optimized [24, 25, 26]. At this point, we do not include quantum Bayesian methods in the package, even though some of the optimization procedure discussed here have their Bayesian counterparts, and may be added in the future updates. The main reason is that, from our numerical experience, the Bayesian approach does not combine well with the tensor-network optimization procedures, and hence cannot be efficiently used to study the performance of protocols in the large number of channel uses, see more discussion in the conclusion section.
@@
The crucial and unique feature of our package is the ability to go to large scale metrological problems, involving multiple-channel uses, whether in adaptive sequential schemes, parallel schemes, or any conceivable protocol structure one may think of, and optimize the protocols without any prior assumption regarding the class of operations allowed, apart from the restriction on the dimensionality of the physical and ancillary spaces involved, see Fig. 1.
@@
As such, our package completely supersedes the TNQMetro package [27], which was designed to deal with optimization of input probes in parallel metrological scheme involving large number of channels. Compared with the previous package, the new one not only allows for a full flexibility of choosing the protocol structure one wants to analyze, but even in the basic parallel channels scenario offers a significantly improved user-friendly interface that allows to deal with large class of problems, including correlated noise models, in a straightforward way.
@@
Interestingly, in case of uncorrelated noise models, it is known that the bounds are tight in the asymptotic limit of large number of channel uses [18, 21]. This fact makes the bounds doubly-powerful. Not only do they serve as a proof that a given protocol is optimal, if it saturates the bound, but also indicate that the protocol is in fact not optimal if one sees a persistent gap between the performance of the numerically found protocol and the bound that does not diminish with the increase of the number of channels used.
@@
The MOP approach has the advantage over ISS that the numerical solution found is guaranteed to be optimal, see Sec. 2.3. Hence, this is the method to be chosen for small-scale problems, where Hilbert spaces dimensions are reasonably small. However, MOP methods cannot be generalized to large-scale problems and combined with tensor-network framework—then, one needs to resort to the ISS approach. The ISS has no guarantee to converge to the true optimal solution, but is constructed in a way that every iteration step cannot reduce the optimized figure-of-merit, see Sec. 2.4.
@@
The main advantage of ISS over MOP is that it naturally combines with tensor-network formalism and may work in the large-$N$ regime. In this case the optimization is performed over each local node in the tensor-network structure separately, whether it represents a part of the input state, control operation or measurement. For this approach one may use the higher-level functions, iss_tnet_parallel_qfi or iss_tnet_adaptive_qfi, dedicated to optimizing standard protocols structures such as parallel (B) or adaptive (C) ones respectively. Alternatively, one can construct a customized structure of the protocol, and use a lower level function iss_opt which performs the iss optimization for arbitrary tensor-network structure, see Sec. 4 for a detailed example.
@@
Finally, the package comes with a number of functions that allow to compute fundamental bounds given just the structure of an individual parameter-encoding quantum channel. The two functions, par_bounds and ad_bounds allow to compute the bounds in case of uncorrelated noise models valid for parallel and adaptive strategies. These bounds are known to be asymptotically saturable for $N\rightarrow\infty$, see Sec. 2.5. The function ad_bounds_correlated, on the other hand, provides bounds in case of correlated noise models, where there is some quantum environment that connects different instances of channels action, and captures the effects of e.g. non-Markovianity. This function provides valid bounds which may be tightened at the expense of numerical complexity, but unlike uncorrelated noise bounds, they are not guaranteed to be tight.
@@
The detailed structure of the package is depicted in Fig. 2. The higher-level optimization functions are collected in the protocols folder, including both MOP and ISS protocols, with the latter one both in the standard as well as tensor-network variant (tnet). This folder also provides a couple of auxiliary functions: a function for computation of QFI of a state (state_qfi), a function for computation of classical Fisher information given a state and measurement (state_cfi) and a function multiple_measurements_qfi that calculates the optimal distribution of $N$ channels into $k$ experiments with $N_{i}$ ($\,N=\sum^{k}_{i=1}N_{i}$) channels. The functions computing bounds are gathered in the bounds folder, where apart from the standard procedure to compute the bounds for finite number of channel uses, there are also asymptotic variants that compute the bounds in the asymptotic limit $N\rightarrow\infty$.
@@
In order to perform the optimization using higher-level functions or compute the bounds, one needs to specify the form of the parameter dependent channel $\Lambda_{\theta}$. The param_channel folder contains all the necessary methods to construct the channel and includes also some exemplary predefined channels in the creators.py file. The channel may be provided in a number of different ways: either through a list of Kraus operatos, or a CJ matrix, or finally as a time integrated quantum Master equation, see Sec. 3.2.
In order to perform optimization of protocols with customized structures one needs to resort to low-level ISS optimization function iss_opt and construct the desired protocol structure using the classes provided in tensors.py file, which allow to specify the elements of the protocol that will be optimized (VarTensor), elements that are fixed (ConstTensor) and elements that include parameter to be estimated (ParamTensor), see Sec. 4.
The package comes also with a number of useful auxiliary functions, combined in qtools.py and other.py, details can be found in the documentation of the package [29]
@@
Sec. 3 is aimed to guide the reader through the high-level optimization functions, going step by step: explaining first how to provide the form of the parameter dependent channel, how to perform the simplest single-channel optimization problems, how to deal with the optimization of parallel and adaptive strategies and finally how to optimize metrological protocols in the most challenging scenarios where noise correlations between subsequent channel uses are present.
@@ 2.1 Quantum channel estimation problem
In a paradigmatic quantum metrological problem the goal is to estimate the value of a single parameter $\theta$ encoded in some physical process described by a quantum channel (formally a Completely Positive Trace Preserving map – CPTP):
@@
where $\mathcal{I}$, $\mathcal{O}$ are Hilbert spaces of input and output quantum systems and $\mathcal{L}\left(\mathcal{H}\right)$ is a set of linear operators on $\mathcal{H}$, that is a super-set of density matrices on $\mathcal{H}$.
@@
where $\mathcal{A}$ is an auxiliary system called ancilla; $\Lambda_{\theta}$ does not act on $\mathcal{A}$, but the entanglement between $\mathcal{I}$ and $\mathcal{A}$ may be sometimes used to enhance the estimation precision in the presence of noise [30, 31, 22, 23]. The resulting output state is
@@
The state $\rho_{\theta}$ is measured by a generalized measurement $\Pi=\{\Pi_{i}\}$ resulting in the probability of outcome $i$ equal to $p_{\theta}(i):=\mathrm{Tr}\left(\rho_{\theta}\Pi_{i}\right)$. The parameter $\theta$ is then estimated using an estimator $\tilde{\theta}(i)$ and the goal is to find the protocol giving results as close as possible to the true value of $\theta$. In case of the simplest single-channel estimation strategy, as depicted in Fig. 1(A), the optimization of the protocol is equivalent to the optimization of the input state $\rho_{0}$ and the measurement $\{\Pi_{i}\}$. However, more complex protocols may additionally involve optimization over control operations $\textrm{C}_{i}$, as in e.g. scenarios (C) and (D) in Fig. 1.
In general, the exact form of the optimal protocol depends on the actual estimation framework chosen, be it frequentist or Bayesian [32, 33, 34], as well as the choice of the cost function/figure of merit to be minimized/maximized. In this work, we adopt the frequentist approach, and focus on optimization of the QFI as the figure of merit to be maximized.
@@ 2.2 Quantum Fisher Information
Given some probabilistic model that specifies probabilities $p_{\theta}(i)$ of obtaining different measurement outcomes $i$ depending on the value of a parameter $\theta$, one can lower bound the achievable mean squared error $\Delta^{2}\tilde{\theta}$ of any locally unbiased estimator $\tilde{\theta}(i)$ via the famous classical Cramér-Rao (CCR) bound [35]:
@@
where $\dot{p}_{\theta}$ denotes the derivative with respect to $\theta$. The bound is asymptotically saturable, in the limit of many repetitions of an experiment $k\rightarrow\infty$, with the help of e.g. maximum-likelihood estimator [35]. Hence, the larger the CFI the better parameter sensitivity of the model.
@@
Moreover, the optimal measurement can in particular be chosen as the projective measurement in the eigenbasis of the SLD.
Combining the above with the CCR bound (4) leads to the quantum Cramér-Rao (QCR) bound, that sets a lower bound on achievable estimation variance when estimating parameter encoded in a quantum state $\rho_{\theta}$:
$\Delta^{2}\tilde{\theta}\geq\frac{1}{kF_{Q}(\rho_{\theta})}.$ (8)
@@
This makes the QFI a fundamental concept in the quantum estimation theory, and makes it meaningful to formulate optimization problems in quantum metrology in a form of optimization of the QFI of the final quantum state obtained at the output of a given metrological protocol.
@@
In case of noiseless unitary estimation models $\Lambda_{\theta}(\rho)=U_{\theta}\rho U_{\theta}^{\dagger}$ the optimal probe states that maximize the QFI are easy to identify analytically [1] and no sophisticated tools are required. In case of noisy models, however, the actual optimization over the protocol is much more involved, as in principle it requires optimization over multi-partite entangled states (B) or multiple control operations $\textrm{C}_{i}$ (C) or both (D).
This package allows for optimization of all kinds of metrological strategies, employing ancillary systems, entanglement and control operations, with the only physical restriction that the operations applied have a definite causal order, and hence can be regarded as a so-called quantum combs in the sense formalized in [36] and further in this text.
@@ 2.3 Minimization over purifications
The MOP method [37, 11, 12, 22] is based on an observation that QFI of a state $\rho_{\theta}\in\mathcal{L}(\mathcal{H})$ is equal to minimization of QFIs over its purifications:
$F_{Q}(\rho_{\theta})=\min_{\ket{\Psi_{\theta}}}F_{Q}(\ket{\Psi_{\theta}}\bra{\Psi_{\theta}}),$ (10)
where |$\ket{\Psi_{\theta}}\in\mathcal{H}\otimes\mathcal{R}$ is a purification of $\rho_{\theta},\,\rho_{\theta}=\mathrm{Tr}_{\mathcal{R}}\ket{\Psi_{\theta}}\bra{\Psi_{\theta}}$ |.
Consider now a parameter encoding channel $\Lambda_{\theta}:\mathcal{L}(\mathcal{I})\mapsto\mathcal{L}(\mathcal{O})$. The idea of minimization over purifications leads to an efficiently computable formula for the channel QFI [37, 11, 12, 22]
$F_{Q}(\Lambda_{\theta})=4\min_{\{K_{k,\theta}\}}\left\|\sum_{k=1}^{r}\dot{K}_{\theta,k}^{\dagger}\dot{K}_{\theta,k}\right\|,$ (11)
@@
This optimization becomes feasible, as one can effectively parametrize derivatives of Kraus operators corresponding to all equivalent Kraus representations as:
@@
where $h\in\mathbb{C}^{r\times r}$ is a Hermitian matrix, while $K_{\theta,k}$ is any fixed Kraus representation of $\Lambda_{\theta}$, e.g. the canonical representation obtained from eigenvectors of the corresponding Choi-Jamiołkowski (CJ) operator [38], see Appendix A for the definition of the CJ operator.
@@
Importantly, this problem may be effectively written as a simple semi-definite program (SDP) [12, 22]
$F_{Q}(\Lambda_{\theta})=4\min_{\begin{subarray}{c}\lambda\in\mathbb{R},\\ h=h^{\dagger}\end{subarray}}\lambda\leavevmode\nobreak\ \leavevmode\nobreak\ \textrm{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ A\succeq 0,$ (16)
@@
$d=\textrm{dim}\mathcal{O}$ and $r$ is the number of Kraus operators. Provided the Hilbert space dimension is small enough, this problem may be solved efficiently using widely available SDP solvers, such as [39]. This optimization is implemented in the mop_channel_qfi function.
Figure 3: Final state resulting from the action of an adaptive protocol on $N$ parameter encoding channels, written via a formal link product operation between the corresponding CJ operators—quantum combs.
Notice, that the dimension of ancilla $\mathcal{A}$ is not included in any of the constraints in (16). This is because MOP gives the maximal QFI without any restriction on ancilla sizes. Furthermore, in order to identify the optimal input state |$\ket{\psi}$ one needs to solve an additional SDP problem, see [18, 22].
MOP can be also used to optimize QFI in case of multiple channel uses. For parallel strategy, Fig. 1(B), one can simply apply the single-channel method, replacing the channel $\Lambda_{\theta}$ with $\Lambda_{\theta}^{(N)}=\Lambda_{\theta}^{\otimes N}$—this is implemented in the mop_parallel_qfi function.
In order to deal with an adaptive strategy, as in Fig. 1(C), one should first use the fact that a concatenated sequence of quantum channels $\mathrm{C}_{i}$ where the free inputs and free output spaces are respectively $\mathcal{H}_{2i}$, $\mathcal{H}_{2i+1}$, $i\in{0,\dots,N-1}$, can be represented as a quantum comb [36]
@@
[…]
The above conditions guarantee that such quantum comb $C$ can always be regarded as a concatenation of quantum channels $C_{i}$, that is $C=C_{0}*...*C_{N-1}$, where * denotes the link product [36], see Appendix A for the definition of the link product.
@@
We may now represent the general adaptive strategy that includes the input state as well as the control operations in a form of a quantum comb
@@
where $\mathit{\Lambda}_{\theta}^{(N)}$ (italics) formally represents the CJ operator of $\Lambda_{\theta}^{(N)}$, see Fig. 3. Here $\mathit{\Lambda}^{(N)}_{\theta}$ does not necessarily need to represent the action of uncorrelated channels as in (20), but may also be a general parameter dependent quantum comb and in this way represent models with correlated noise or non-local character of estimated parameter.
Optimization over adaptive protocols is thus equivalent to maximization of the final QFI over quantum combs $C$. Using again the MOP idea, this problem may be effectively written down as an SDP (this time slightly more involved) of the following form [19, 20, 22]:
$F_{\textrm{Q}}^{(N)}\left(\Lambda_{\theta}\right)=4\min_{\begin{subarray}{c}\lambda\in\mathbb{R},Q^{(k)}\\ h=h^{\dagger}\end{subarray}}\lambda\leavevmode\nobreak\ \leavevmode\nobreak\ \textrm{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ A\succeq 0,$ (23)
@@
$\left(\begin{array}[]{c|c}\mathbbm{1}_{\mathcal{I}_{N-1}}\otimes Q^{(N-1)}&\ket{\dot{\tilde{K}}_{1,1}(h)}\ldots\ket{\dot{\tilde{K}}_{r,d}(h)}\\ \hline\cr\bra{\dot{\tilde{K}}_{1,1}(h)}&\\ \vdots&\lambda\mathbb{1}_{dr}\\ \bra{\dot{\tilde{K}}_{r,d}(h)}&\end{array}\right)$ (24)
@@
|$\ket{\dot{{K}}_{i,k}^{(N)}(h)}\in\mathcal{O}_{0}\otimes\mathcal{I}_{0}\otimes\dots\otimes\mathcal{O}_{N-1}\otimes\mathcal{I}_{N-1},$ (26)

### v2

@@ Abstract
QMetro++ is a Python package that provides a set of tools for identifying optimal estimation protocols that maximize quantum Fisher information (QFI). Optimization can be performed for arbitrary configurations of input states, parameter-encoding channels, noise correlations, control operations, and measurements. The use of tensor networks and an iterative see-saw algorithm allows for an efficient optimization even in the regime of a large number of channel uses ($N\approx 100$). Additionally, the package includes implementations of the recently developed methods for computing fundamental upper bounds on QFI, which serve as benchmarks for assessing the optimality of numerical optimization results. All functionalities are wrapped up in a user-friendly interface which enables the definition of strategies at various levels of detail.
@@ 1 Introduction
(B) parallel small NN mop_parallel_qfi, iss_parallel_qfi 3.3
large NN iss_tnet_parallel_qfi
@@
small NN mop_adaptive_qfi, iss_adaptive_qfi 3.4 3.5
(C) large NN iss_tnet_adaptive_qfi
@@
Figure 2: File structure of the package with a selection of the most important functions. Classes from tensors.py file are displayed in more detail in Fig. 10. All functions, classes and methods are listed and described in the package documentation [19].
@@
Quantum metrology [24, 46, 41, 42] is now rightfully regarded as one of the main pillars of the rapidly developing field of quantum technologies [45]. It focuses on the ways to exploit delicate quantum features of light and matter to enhance the sensitivity of practical measuring devices. Apart from numerous experimental breakthroughs in the field, including the use of squeezed states of light in gravitational wave detectors [43, 23], quantum-enhanced magnetometry [48], gravitometry [9], atomic clocks [40] and many others, the field has also enjoyed significant theoretical progress with numerous analytical and numerical methods being developed that help to design better metrological protocols as well as understand the fundamental limitations of quantum-enhanced protocols in the presence of decoherence [20, 16, 17, 15, 13, 51, 11, 50, 2, 33, 32, 31, 34].
@@
The purpose of this package is to make these advanced and powerful numerical methods accessible to the community. We realize that despite the fact that the core elements of the methods have been discussed in the literature already for a couple of years, they are still not widely used in the field due to their somewhat steep ’initial learning curve’ and the fact that there is no single publication that provides a unified presentation of all of them, not to mention a numerical package. Moreover, we significantly expand the applicability of the latest tensor network-based methods to include general protocol structures, which allows for dealing with e.g. collisional metrological schemes.
@@
Not so rarely, we have come across research conducted in the field, where the problem of optimization of quantum metrological protocols has been approached using general-purpose optimization tools (including even the latest AI methods), which in the end did not result in a particularly efficient numerical algorithm, the results of numerical optimization had no optimality guarantee and were not benchmarked against known fundamental precision bounds [citations intentionally hidden here!].
@@
This package adopts the frequentist approach to quantum metrology and focuses on quantum Fisher information (QFI) as the figure-of-merit to be optimized [27, 7, 39]. At this point, we do not include quantum Bayesian methods in the package, even though some of the optimization procedures discussed here have their Bayesian counterparts, and may be added in the future updates. The main reason is that, from our numerical experience, the Bayesian approach does not combine well with the tensor-network optimization procedures, and hence cannot be efficiently used to study the performance of protocols in a large number of channel uses, see further discussion in the conclusion section.
@@
The crucial and unique feature of our package is the ability to go to large-scale metrological problems, involving multiple channel uses, whether in adaptive sequential schemes, parallel schemes, or any conceivable protocol structure one may think of, and optimize the protocols without any prior assumption regarding the class of operations allowed, apart from the restriction on the dimensionality of the physical and ancillary spaces involved, see Fig. 1.
@@
As such, our package completely supersedes the TNQMetro package [10], which was designed to deal with optimization of input probes in a parallel metrological scheme involving a large number of channels. Compared with the previous package, the new one not only allows for a full flexibility of choosing the protocol structure one wants to analyze, but even in the basic parallel channels scenario offers a significantly improved user-friendly interface that allows to deal with a large class of problems, including correlated noise models, in a straightforward way.
@@
Interestingly, in case of uncorrelated noise models, it is known that the bounds are tight in the asymptotic limit of large number of channel uses [50, 32]. This fact makes the bounds powerful in two ways. Not only do they serve as a proof that a given protocol is optimal, if it saturates the bound, but also indicate that the protocol is in fact not optimal if one sees a persistent gap between the performance of the numerically found protocol and the bound that does not diminish with the increase of the number of channels used.
@@
The MOP approach has the advantage over ISS that the numerical solution found is guaranteed to be optimal, see Sec. 2.3. Hence, this is the method to be chosen for small-scale problems, where dimensions of the Hilbert spaces are reasonably small. However, MOP methods cannot be generalized to large-scale problems and combined with tensor-network framework—then, one needs to resort to the ISS approach. The ISS has no guarantee to converge to the true optimal solution, but is constructed in a way that every iteration step cannot reduce the optimized figure-of-merit, see Sec. 2.4.
@@
The main advantage of ISS over MOP is that it naturally combines with tensor-network formalism and may work in the large-$N$ regime. In this case the optimization is performed over each local node in the tensor-network structure separately, whether it represents a part of the input state, control operation or measurement. For this approach one may use the higher-level functions, iss_tnet_parallel_qfi or iss_tnet_adaptive_qfi, dedicated to optimizing standard protocol structures such as parallel (B) or adaptive (C) ones respectively. Alternatively, one can construct a customized structure of the protocol, and use a lower-level function iss_opt which performs the iss optimization for arbitrary tensor-network structure, see Sec. 4 for a detailed example.
@@
Finally, the package comes with a number of functions that allows one to compute fundamental bounds given just the structure of an individual parameter-encoding quantum channel. The two functions, par_bounds and ad_bounds allow to compute the bounds in case of uncorrelated noise models valid for parallel and adaptive strategies. These bounds are known to be asymptotically saturable for $N\rightarrow\infty$, see Sec. 2.5. The function ad_bounds_correlated, on the other hand, provides bounds in case of correlated noise models, where there is some quantum environment that connects different instances of channels action, and captures the effects of e.g. non-Markovianity. This function provides valid bounds which may be tightened at the expense of numerical complexity, but unlike uncorrelated noise bounds, they are not guaranteed to be tight.
@@
The detailed structure of the package is depicted in Fig. 2. The higher-level optimization functions are collected in the protocols folder, including both MOP and ISS protocols, with the latter one in the standard as well as the tensor-network variant (tnet). This folder also provides a couple of auxiliary functions: a function for computation of QFI of a state (state_qfi), a function for computation of classical Fisher information given a state and measurement (state_cfi) and a function multiple_measurements_qfi that calculates the optimal distribution of $N$ channels into $k$ experiments with $N_{i}$ ($N=\sum^{k}_{i=1}N_{i}$) channels. The functions computing bounds are gathered in the bounds folder, where apart from the standard procedure to compute the bounds for finite number of channel uses, there are also asymptotic variants that compute the bounds in the asymptotic limit $N\rightarrow\infty$.
@@
In order to perform the optimization using higher-level functions or compute the bounds, one needs to specify the form of the parameter-dependent channel $\Lambda_{\theta}$. The param_channel folder contains all the necessary methods to construct the channel and includes also some exemplary predefined channels in the creators.py file. The channel may be provided in a number of different ways: either through a list of Kraus operators, or a CJ matrix, or finally as a time integrated quantum Master equation, see Sec. 3.2.
For optimization of protocols with customized structures one needs to resort to low-level ISS optimization function iss_opt and construct the desired protocol structure using the classes provided in tensors.py file, which allow to specify the elements of the protocol that will be optimized (VarTensor), elements that are fixed (ConstTensor) and elements that include parameter to be estimated (ParamTensor), see Sec. 4.
The package comes also with a number of useful auxiliary functions, combined in qtools.py module. Among them functions that allow to switch between different representations of quantum channels like Kraus, Choi-Jamiołkowski and Lindbladian operators. We describe them throughout Sec. 3 and 4.
@@
Sec. 3 aims to guide the reader through the high-level optimization functions, going step by step: explaining first how to provide the form of the parameter-dependent channel, how to perform the simplest single-channel optimization problems, how to deal with the optimization of parallel and adaptive strategies and finally how to optimize metrological protocols in the most challenging scenarios where noise correlations between subsequent channel uses are present.
@@ 2.1 Quantum channel estimation problem
In a paradigmatic quantum metrological problem, the goal is to estimate the value of a single parameter $\theta$ encoded in some physical process described by a quantum channel (formally a Completely Positive Trace-Preserving map – CPTP):
@@
where $\mathcal{I}$, $\mathcal{O}$ are Hilbert spaces of input and output quantum systems and $\mathcal{L}\left(\mathcal{H}\right)$ is a set of linear operators on $\mathcal{H}$, that is a superset of density matrices on $\mathcal{H}$.
@@
where $\mathcal{A}$ is an auxiliary system called an ancilla; $\Lambda_{\theta}$ does not act on $\mathcal{A}$, but the entanglement between $\mathcal{I}$ and $\mathcal{A}$ may sometimes be used to enhance the estimation precision in the presence of noise [22, 29, 31, 34]. The resulting output state is
@@
The state $\rho_{\theta}$ is measured by a generalized measurement $\Pi=\{\Pi_{i}\}$ resulting in the probability of outcome $i$ equal to $p_{\theta}(i):=\mathrm{Tr}\left(\rho_{\theta}\Pi_{i}\right)$. The parameter $\theta$ is then estimated using an estimator $\tilde{\theta}(i)$ and the goal is to find the protocol producing results as close as possible to the true value of $\theta$. In case of the simplest single-channel estimation strategy, as depicted in Fig. 1(A), the optimization of the protocol is equivalent to optimizing the input state $\rho_{0}$ and the measurement $\{\Pi_{i}\}$. However, more complex protocols may additionally involve optimization over control operations $\textrm{C}_{i}$, as in scenarios (C) and (D) in Fig. 1.
In general, the exact form of the optimal protocol depends on the actual estimation framework chosen, be it frequentist or Bayesian [26, 38, 14], as well as the choice of the cost function/figure of merit to be minimized/maximized. In this work, we adopt the frequentist approach, and focus on optimizing the QFI as the figure of merit to be maximized.
@@ 2.2 Quantum Fisher Information
Given some probabilistic model that specifies probabilities $p_{\theta}(i)$ of obtaining different measurement outcomes $i$ depending on the value of a parameter $\theta$, one can lower-bound the achievable mean squared error $\Delta^{2}\tilde{\theta}$ of any locally unbiased estimator $\tilde{\theta}(i)$ via the famous classical Cramér-Rao (CCR) bound [28]:
@@
where $\dot{p}_{\theta}$ denotes the derivative with respect to $\theta$. The bound is asymptotically saturable, in the limit of many repetitions of an experiment $k\rightarrow\infty$, with the help of e.g. a maximum-likelihood estimator [28]. Hence, the larger the CFI, the better parameter sensitivity of the model.
@@
Moreover, the optimal measurement can, in particular, be chosen as the projective measurement in the eigenbasis of the SLD.
Combining the above with the CCR bound (4) leads to the quantum Cramér-Rao (QCR) bound that sets a lower bound on achievable estimation variance when estimating a parameter encoded in a quantum state $\rho_{\theta}$:
$\Delta^{2}\tilde{\theta}\geq\frac{1}{kF_{Q}(\rho_{\theta})},$ (8)
@@
This makes the QFI a fundamental concept in quantum estimation theory, and makes it meaningful to formulate optimization problems in quantum metrology in the form of an optimization of the QFI of the final quantum state obtained at the output of a given metrological protocol.
@@
In case of noiseless unitary estimation models $\Lambda_{\theta}(\rho)=U_{\theta}\rho U_{\theta}^{\dagger}$ the optimal probe states that maximize the QFI are easy to identify analytically [24] and no sophisticated tools are required. In case of noisy models, however, the actual optimization over the protocol is much more involved, as in principle it requires optimization over multipartite entangled states (B) or multiple control operations $\textrm{C}_{i}$ (C) or both (D).
This package allows for optimization of all kinds of metrological strategies, employing ancillary systems, entanglement and control operations, with the only physical restriction that the operations applied have a definite causal order, and hence can be regarded as so-called quantum combs in the sense formalized in [12] and further in this text.
@@ 2.3 Minimization over purifications
The MOP method [21, 20, 16, 31] is based on an observation that the QFI of a state $\rho_{\theta}\in\mathcal{L}(\mathcal{H})$ is equal to the minimum of QFI over its purifications:
@@
[…]
Consider now a parameter-encoding channel $\Lambda_{\theta}:\mathcal{L}(\mathcal{I})\mapsto\mathcal{L}(\mathcal{O})$. The idea of minimization over purifications leads to an efficiently computable formula for the channel QFI [21, 20, 16, 31]
$F_{Q}(\Lambda_{\theta})=4\min_{\{K_{\theta,k}\}}\left\|\sum_{k=1}^{r}\dot{K}_{\theta,k}^{\dagger}\dot{K}_{\theta,k}\right\|,$ (11)
where $\|\cdot\|$ is the operator norm and the minimization is performed over all equivalent Kraus representations of the channel, such that:
@@
This optimization becomes feasible, as one can effectively parametrize the derivatives of the Kraus operators corresponding to all equivalent Kraus representations as:
@@
where $h\in\mathbb{C}^{r\times r}$ is a Hermitian matrix and $K_{\theta,k}$ is any fixed Kraus representation of $\Lambda_{\theta}$, e.g. the canonical representation obtained from eigenvectors of the corresponding Choi-Jamiołkowski (CJ) operator [5], see Appendix A for the definition of the CJ operator.
@@
Importantly, this problem may be effectively written as a simple semidefinite program (SDP) [16, 31]
$F_{Q}(\Lambda_{\theta})=4\min_{\begin{subarray}{c}\lambda\in\mathbb{R},\\ h=h^{\dagger}\end{subarray}}\lambda\penalty 10000\ \penalty 10000\ \textrm{s.t.}\penalty 10000\ \penalty 10000\ A\succeq 0,$ (16)
@@
$d=\textrm{dim}\mathcal{O}$ and $r$ is the number of Kraus operators. Provided the dimension of the relevant Hilbert space is small enough, this problem may be solved efficiently using widely available SDP solvers, such as [3]. This optimization is implemented in the mop_channel_qfi function.
Figure 3: Final state resulting from the action of an adaptive protocol on NN parameter encoding channels, written via a formal link product operation between the corresponding CJ operators—quantum combs.
[…]
MOP can also be used to optimize QFI in case of multiple channel uses. For the parallel strategy, Fig. 1(B), one can simply apply the single-channel method, replacing the channel $\Lambda_{\theta}$ with $\Lambda_{\theta}^{(N)}=\Lambda_{\theta}^{\otimes N}$—this is implemented in the mop_parallel_qfi function.
In order to address an adaptive strategy, as in Fig. 1(C), one should first use the fact that a concatenated sequence of quantum channels $\mathrm{C}_{i}$ where the free inputs and free output spaces are respectively $\mathcal{H}_{2i}$, $\mathcal{H}_{2i+1}$, $i\in{0,\dots,N-1}$, can be represented as a quantum comb [12]
@@
$\mathrm{Tr}_{2k-1}C^{(k)}=\text{1}_{2k-2}\otimes C^{(k-1)},\quad k\in\{1,...,N\}.$
The above conditions guarantee that such quantum comb $C$ can always be regarded as a concatenation of quantum channels $C_{i}$, namely $C=C_{0}*...*C_{N-1}$, where * denotes the link product [12], see Appendix A for the definition of the link product.
@@
We may now represent the general adaptive strategy that includes the input state as well as the control operations in the form of a quantum comb
@@
where $\mathit{\Lambda}_{\theta}^{(N)}$ (italics) formally represents the CJ operator of $\Lambda_{\theta}^{(N)}$, see Fig. 3. Here $\mathit{\Lambda}^{(N)}_{\theta}$ does not necessarily need to represent the action of uncorrelated channels as in (20), but may also be a general parameter-dependent quantum comb and in this way represent models with correlated noise or non-local character of estimated parameter.
Optimization over adaptive protocols is thus equivalent to maximizing the final QFI over quantum combs $C$. Using again the MOP idea, this problem may be effectively written down as an SDP (this time slightly more involved) of the following form [2, 33, 31]:
$F_{\textrm{Q}}^{(N)}\left(\Lambda_{\theta}\right)=4\min_{\begin{subarray}{c}\lambda\in\mathbb{R},Q^{(k)}\\ h=h^{\dagger}\end{subarray}}\lambda\penalty 10000\ \penalty 10000\ \textrm{s.t.}\penalty 10000\ \penalty 10000\ A\succeq 0,$ (23)
@@
[…]
@@
[…]